OpenStudy (anonymous):

\[t^2 y''-4 t y'+6y==0\]

OpenStudy (anonymous):

\[y_1=t^2\]

OpenStudy (anonymous):

\[y_2= v(t) y_1\] \[y_2= v(t) t^2\] \[y'=v' t^2+v 2t\] \[y''=4v' t+v'' t^2+ 2v\] plugging into differential equation, we get \[v''(t) t^4=0\] \[v'(t) =D\] \[v(t) =Dt+C\] \[y_2= v t^2\] \[y_2 =t^3\]